B. $3,680.90 C. $6,884.70 D. $3,690.38 E. $7,021.38 M 3. What is the present value of $13,100 to be received 3 years from today if the discount rate is 5 percent? A. $10,767.25 B. $11,231.14 C.$7,860.00 D. S11,316.27 E.$11,882.09 M 4. Three years ago, you invested $2,900.00. Today, it is worth $3,700.00. What rate of interest did you earn? A. 4.23 percent B. 8.46 percent C. 4.63 percent D. 0.71 percent E. 3.63 percent M5. Bob bought some land costing $16,340. Today, that same land is valued at $46,717. How long has Bob owned this land if the price of land has been increasing at 4 percent per year? A. 27.85 years B. 26.78 years ?. 11.00 years D. 28.96 years E. 11.44 years Page 2c
Answer #1

3. Computation of Present Value

Present Value = Future Value / (1+r)^3

Present Value = $13100 / (1.05)^3

Present Value =  $13100 * 0.8638

Present Value = $11316.27

4. Computation Rate of Interest

Future Value = Present Value * (1+r)^n

$3700 = $2900 * (1+r)^3

Cube root (1.2759) = 1+r

1.0846 = 1+r

Rate of Interest = 8.46%

5. Computation of Number of years

Future Value = Present Value * (1+r)^n

$46717 = $16340 * (1.04)^n

2.8591 = 1.04^n

1.04^26.78 = 1.04^n

Number of Years = 26.78 Years